The directions on a bag of cement are to mix 1 part cement, 2 parts sand, and 3 part gravel. If you have 14 pounds of sand, how many pounds of cement and how many pounds of gravel should be mixed with this sand?
Found 2 solutions by richwmiller, ptaylor:
Answer by richwmiller(17219)

14=2*7
7 pounds of cement
21 pounds of gravel
Are the instructions talking about weight or volume?

Answer by ptaylor(2198)

OK. There is 6 parts to this mixture: 1 part cement(1/6), 2 parts sand(2/6) and 3 parts gravel(3/6).
Let x=total amount of mixture
(1/6)x= -------cement
2/6=(1/3)x=-----sand
3/6=(1/2)x=------gravel
Now we are told that:
(1/3)x=14 or
x=42 pounds----------------------total amount of mixture
(1/2)x=(1/2)*42=21 pounds----amount of gravel
CK
(1/6)*42=7
(1/3)*42=14
(1/2)*42=21
21+14+7=42
42=42

Hope this helps---ptaylor